Verify that the equations i) z =sqrt (2x + a )+ sqrt(2y + b) and ii)z^2+u=2(1+l ^x)(x+ly) are both complete integrals of the PDEz=1/p+1/q . Also show that the complete integral (ii) is the envelope of one parameter sub-system obtained by taking b=-a/l -μ/1+l in the solution (i)
Expert's answer
1
Expert's answer
2020-03-23T11:50:34-0400

i) Given

"\ \ \ z=\sqrt{2x+a}+\sqrt{2y+b} \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (1)"

Differentiate (1) partially with respect to "x"


"\\ \ \ {\partial z \over \partial x}={2 \over 2\sqrt{2x+a}}={ 1\over \sqrt{2x+a}} \ \ \ \ \ \ \ \ \ \ \ (2)"

Differentiate (1) partially with respect to "y"


"\\ \ \ {\partial z \over \partial y}={2 \over 2\sqrt{2+b}}={ 1\over \sqrt{2y+b}} \ \ \ \ \ \ \ \ \ \ \ (3)"

From (2) and (3)


"p=\sqrt{2x+a},\ q=\sqrt{2y+b}"

Substitutiing in (1) we get


"z={1 \over p}+{1 \over q}"

ii) Given


"\ \ z^2+\mu=2(1+\lambda^{-1})(x+\lambda y) \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (4)"

Differentiate (4) partially with respect to "x"


"2z{\partial z \over \partial x}=2(1+\lambda^{-1})"


"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ z{\partial z \over \partial x}=1+\lambda^{-1} \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (5)"

Differentiate (4) partially with respect to "y"


"2z{\partial z \over \partial y}=2\lambda(1+\lambda^{-1})"

"\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ z{\partial z \over \partial y}=\lambda(1+\lambda^{-1}) \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ (6)"

From (5) and (6)


"p={1 \over z}(1+\lambda^{-1}),\ q={1 \over z}\lambda(1+\lambda^{-1})"

Then

"{1 \over p}+{1 \over q}={z \over 1+\lambda^{-1}}+{z \over \lambda(1+\lambda^{-1})}"

"{1 \over p}+{1 \over q}=z{\lambda+1 \over \lambda(1+\lambda^{-1})}"

"{1 \over p}+{1 \over q}=z{\lambda+1 \over \lambda+1}"

Hence


"z={1 \over p}+{1 \over q}"

 Show that the complete integral (ii) is the envelope of one parameter sub-system obtained by taking 


"b=-{a \over \lambda}-{\mu \over1+\lambda}"

Given


"z=\sqrt{2x+a}+\sqrt{2y+b}"

Then


"f(x, y, z, a, b)=\sqrt{2x+a}+\sqrt{2y+b}-z=0"

"{\partial f \over \partial a}={1 \over 2\sqrt{2x+a}}+{1 \over 2\sqrt{2y+b}}\cdot{db \over da}=0"

"\sqrt{2y+b}=-{db \over da}\sqrt{2x+a}"

Let

"b=-{a \over \lambda}+c"

Then


"\sqrt{2y+b}={1 \over \lambda}\sqrt{2x+a}"

"2y+b={2 \over \lambda^2}x+{1 \over \lambda^2}a"

"z=(1+{1 \over \lambda})\sqrt{2x+a}"

"z^2={(1+\lambda)^2 \over \lambda^2}(2x+a)"

If


"z^2+\mu=2(1+\lambda^{-1})(x+\lambda y)"

Then


"{(1+\lambda)^2 \over \lambda^2}(2x+a)={2(1+\lambda) \over \lambda}(x+\lambda y)-\mu"

"{1+\lambda \over \lambda}(2x+a)=2x+\lambda(2y)-{\lambda \over 1+\lambda}\mu"

"2(1+\lambda)x+(1+\lambda)a=2\lambda x+\lambda^2({2 \over \lambda^2}x+{1 \over \lambda^2}a-b)-{\lambda^2 \over 1+\lambda}\mu"

"(1+\lambda)a=a-\lambda^2b-{\lambda^2 \over 1+\lambda}\mu"

"b=-{a \over \lambda}-{\mu \over1+\lambda}"

The complete integral (ii) is the envelope of one parameter sub-system obtained by taking 


"b=-{a \over \lambda}-{\mu \over1+\lambda}"